# JaxMARL: Multi-Agent RL Environments and Algorithms in JAX

### v4

@@ A.3. Spatial-Temporal Representations of Matrix Games (STORM)
MeltingPot is implemented in Lua (Ierusalimschy, 2006) where as ours is a vectorized implementation in Jax.
@@
We deem the coin shuffling especially crucial because even large environments representing POMDPs, such as SMAC, can be solved without the need for memory if they lack sufficient randomness (Ellis et al., 2022).
A.4. Hanabi
There are a few details that differ between our Hanabi implementation and the original Hanabi Learning Environment (HLE). The most notable of these is how we choose to represent card knowledge information in the agents’ observation. In the HLE, card knowledge is observed as a colour/rank if there has been an explicit hint about a given card. As a separate feature, implicit card knowledge is represented as possible colours/ranks if there has not been an explicit hint that indicates a given card is not that colour/rank. We, on the other hand, combine implicit and explicit card knowledge, by only maintaining a representation of implicit card knowledge, which reduces to explicit card knowledge in the event an explicit hint is given about a card. This is because all possible colours/ranks are represented as 1s, whilst all ruled out colours/ranks are represented as 0s. By giving an explicit hint, all but one colour/rank are ruled out, leaving a one-hot encoding of the explicit card knowledge. We implement card knowledge this way, because knowledge updates are implemented via tensor calculus using JAX Numpy arrays of fixed shape and data type.
Appendix B Value-Based MARL Methods and Implementation details
@@
Key features of our framework include parameter sharing, a recurrent neural network (RNN) for agents, an epsilon-greedy exploration strategy with linear decay, a uniform experience replay buffer, and the incorporation of Double Deep -Learning (DDQN) Van Hasselt et al. (2016) techniques to enhance training stability.
@@
IQL (Independent -Learners) is a straightforward adaptation of Deep -Learning to multi-agent scenarios. It features multiple -Learner agents that operate independently, optimizing their individual returns. This approach follows a decentralized learning and decentralized execution pipeline.
VDN (Value Decomposition Networks) extends -Learning to multi-agent scenarios with a centralized-learning-decentralized-execution framework. Individual agents approximate their own action’s -Value, which is then summed during training to compute a jointed for the global state-action pair. Back-propagation of the global DDQN loss in respect to a global team reward optimizes the factorization of the jointed -Value.
@@
Appendix C Training Results
C.1. MABrax
@@
The performance of IPPO on HalfCheeta and Walker is reported in Figure 9, with hyperparameters reported in Table 4.
C.2. MPE
@@
Performance of -Learning baselines in all the MPE scenarios are reported in Figure 10. The upper row represents cooperative scenarios, with results for all our -learning baselines reported. The bottom row refers to competitive scenarios, and results for IQL are divided by agent types. Hyperparameters are given in Table 10
C.3. SMAX
The performance of IPPO in SMAX versus MAPPO in SMAC is shown in Figure 11 while the performance of our -learning baselines is reported in Figure 12. We do not report them together because their hyperparameters were tuned over a different number of timesteps. Hyperparameters for IPPO and the -learning methods are given in Tables 6 and 10 respectively.
Appendix D Hyperparameters
| Value | Ant | HalfCheetah | Walker |
| VF_COEF | 4.5 | 0.14 | 1.9 |
| ENT_COEF | |||
| LR | |||
| NUM_ENVS | 64 | – | – |
| NUM_STEPS | 300 | – | – |
| TOTAL_TIMESTEPS | – | – | |
| NUM_MINIBATCHES | 4 | – | – |
| GAMMA | 0.99 | – | – |
| GAE_LAMBDA | 1.0 | – | – |
| CLIP_EPS | 0.2 | – | – |
| MAX_GRAD_NORM | 0.5 | – | – |
| ACTIVATION | tanh | – | – |
| ANNEAL_LR | True | – | – |
| Hyperparameter | Value |
| LR | |
| NUM_ENVS | |
| NUM_STEPS | |
| TOTAL_TIMESTEPS | |
| UPDATE_EPOCHS | |
| NUM_MINIBATCHES | |
| GAMMA | |
| GAE_LAMBDA | |
| CLIP_EPS | |
| ENT_COEF | |
| VF_COEF | |
| MAX_GRAD_NORM | |
| ACTIVATION | tanh |
| ANNEAL_LR | True |
| Hyperparameter | Value |
| LR | |
| NUM_ENVS | |
| NUM_STEPS | |
| TOTAL_TIMESTEPS | |
| UPDATE_EPOCHS | |
| NUM_MINIBATCHES | |
| GAMMA | |
| GAE_LAMBDA | |
| CLIP_EPS | |
| SCALE_CLIP_EPS | False |
| ENT_COEF | |
| VF_COEF | |
| MAX_GRAD_NORM | |
| ACTIVATION | relu |
| Hyperparameter | Value |
| LR | |
| NUM_ENVS | |
| NUM_STEPS | |
| TOTAL_TIMESTEPS | |
| UPDATE_EPOCHS | |
| NUM_MINIBATCHES | |
| GAMMA | |
| GAE_LAMBDA | |
| CLIP_EPS | |
| ENT_COEF | |
| VF_COEF | |
| MAX_GRAD_NORM | |
| ACTIVATION | relu |
| ANNEAL_LR | True |
| NUM_FC_LAYERS | 2 |
| LAYER_WIDTH | 512 |
| Hyperparameter | Value |
| LR | |
| NUM_ENVS | |
| NUM_STEPS | |
| TOTAL_TIMESTEPS | |
| UPDATE_EPOCHS | |
| NUM_MINIBATCHES | |
| GAMMA | |
| GAE_LAMBDA | |
| CLIP_EPS | |
| ENT_COEF | |
| VF_COEF | |
| MAX_GRAD_NORM | |
| ACTIVATION | tanh |
| ANNEAL_LR | True |
| NUM_EVALS |
| Hyperparameter | Value |
| NUM_ENVS | |
| NUM_STEPS | |
| BUFFER_SIZE | |
| BUFFER_BATCH_SIZE | |
| TOTAL_TIMESTEPS | |
| AGENT_HIDDEN_DIM | |
| AGENT_INIT_SCALE | |
| EPSILON_START | |
| EPSILON_FINISH | |
| EPSILON_ANNEAL_TIME | |
| MIXER_EMBEDDING_DIM* | |
| MIXER_HYPERNET_HIDDEN_DIM* | |
| MIXER_INIT_SCALE* | |
| MAX_GRAD_NORM | |
| TARGET_UPDATE_INTERVAL | |
| LR | |
| EPS_ADAM | |
| WEIGHT_DECAY_ADAM | |
| GAMMA | |
| NUM_TEST_EPISODES | |
| TEST_INTERVAL |
(* Parameters specific to QMix.)
| Hyperparameter | Value |
| NUM_ENVS | |
| NUM_STEPS | |
| BUFFER_SIZE | |
| BUFFER_BATCH_SIZE | |
| TOTAL_TIMESTEPS | |
| AGENT_HIDDEN_DIM | |
| AGENT_INIT_SCALE | |
| EPSILON_START | |
| EPSILON_FINISH | |
| EPSILON_ANNEAL_TIME | |
| MIXER_EMBEDDING_DIM* | |
| MIXER_HYPERNET_HIDDEN_DIM* | |
| MIXER_INIT_SCALE* | |
| MAX_GRAD_NORM | |
| TARGET_UPDATE_INTERVAL | |
| LR | |
| EPS_ADAM | |
| WEIGHT_DECAY_ADAM | |
| GAMMA | |
| NUM_TEST_EPISODES | |
| TEST_INTERVAL |
(* Parameters specific to QMix.)

### v5

@@ B.2 Spatial-Temporal Representations of Matrix Games (STORM)
MeltingPot is implemented in Lua [27] where as ours is a vectorized implementation in JAX.
@@
We deem the coin shuffling especially crucial because even large environments representing POMDPs, such as SMAC, can be solved without the need for memory if they lack sufficient randomness [15].
B.3 Coin Game
Coin Game is a two-player grid-world environment which emulates social dilemmas such as the iterated prisoner’s dilemma [57]. Used as a benchmark for the general-sum setting, it expands on simpler social dilemmas by adding a high-dimensional state. Two players, ‘red’ and ‘blue’ move in a grid world and are each awarded 1 point for collecting any coin. However, ‘red’ loses 2 points if ‘blue’ collects a red coin and vice versa. Thus, if both agents ignore colour when collecting coins their expected reward is 0.
B.4 Switch Riddle
Originally used to illustrate the Differentiable Inter-Agent Learning algorithm [16], Switch Riddle is a simple cooperative communication environment that we include as a debugging tool. prisoners held by a warden can secure their release by collectively ensuring that each has passed through a room with a light bulb and a switch. Each day, a prisoner is chosen at random to enter this room. They have three choices: do nothing, signal to the next prisoner by toggling the light, or inform the warden they think all prisoners have been in the room. The game ends when a prisoner informs the warden or the maximum time steps are reached. The rewards are +1 if the prisoner informs the warden, and all prisoners have been in the room, -1 if the prisoner informs the warden before all prisoners have taken their turn, and 0 otherwise, including when the maximum time steps are reached. We benchmark using the implementation from [68].
B.5 Hanabi
Hanabi is a fully cooperative partially observable multiplayer card game, where players can observe other players’ cards but not their own. To win, the team must play a series of cards in a specific order while sharing only a limited amount of information between players. As reasoning about the beliefs and intentions of other agents is central to performance, it is a common benchmark for ZSC and ad-hoc teamplay research. Our implementation is inspired by the Hanabi Learning Environment [3] and includes custom configurations for varying game settings, such as the number of colours/ranks, number of players, and number of hint tokens. Compared to the Hanabi Learning Environment, which is written in C++ and split over dozens of files, our implementation is a single easy-to-read Python file, which simplifies interfacing with the library and running experiments.
B.6 Overcooked
Inspired by the popular videogame of the same name, Overcooked is commonly used for assessing fully cooperative and fully observable Human-AI task performance. The aim is to quickly prepare and deliver soup, which involves putting three onions in a pot, cooking the soup, and serving it into bowls. Two agents, or cooks, must coordinate to effectively divide the tasks to maximise their common reward signal. Our implementation mimics the original from Overcooked-AI [10], including all five original layouts and a simple method for creating additional ones. For a discussion on the limitations of the Overcooked-AI environment, see [33].
B.7 Multi-Agent Particle Environments (MPE)
The multi-agent particle environments feature a 2D world with simple physics where particle agents can move, communicate, and interact with fixed landmarks. Each specific environment varies the format of the world and the agents’ abilities, creating a diverse set of tasks that include both competitive and cooperative settings. We implement all the MPE scenarios featured in the PettingZoo library and the transitions of our implementation map exactly to theirs. We additionally include a fully cooperative predator-prey variant of simple tag, presented in [49]. The code is structured to allow for straightforward extensions, enabling further tasks to be added.
B.8 Multi-Agent Brax (MABrax)
MABrax is a derivative of Multi-Agent MuJoCo [49], an extension of the MuJoCo Gym environment [62] that is commonly used for benchmarking continuous multi-agent robotic control.
Our implementation utilises Brax[19] as the underlying physics engine and includes five of Multi-Agent MuJoCo’s multi-agent factorisation tasks, where each agent controls a subset of the joints and only observes the local state.
The task descriptions mirror those from Gymnasium-Robotics [13].
Appendix C JaxMARL’s API
The interface of JaxMARL is inspired by PettingZoo [61] and Gymnax. We designed it to be a simple and easy-to-use interface for a wide-range of MARL problems. An example of instantiating an environment from JaxMARL’s registry and executing one transition is presented in Figure 10.
This key is separated from the internal state for clarity.
Thus, we do not use PettingZoo’s agent iterator.
Appendix D Value-Based MARL Methods and Implementation details
@@
Key features of our framework include parameter sharing, a recurrent neural network (RNN) for agents, an epsilon-greedy exploration strategy with linear decay, a uniform experience replay buffer, and the incorporation of Double Deep Q-Learning (DDQN) [64] techniques to enhance training stability. We stored the replay buffer in GPU memory using Flashbax [63].
@@
IQL (Independent Q-Learners) is a straightforward adaptation of Deep Q-Learning to multi-agent scenarios. It features multiple Q-Learner agents that operate independently, optimizing their individual returns. This approach follows a decentralized learning and decentralized execution pipeline.
VDN (Value Decomposition Networks) extends Q-Learning to multi-agent scenarios with a centralized-learning-decentralized-execution framework. Individual agents approximate their own action’s -Value, which is then summed during training to compute a jointed for the global state-action pair. Back-propagation of the global DDQN loss in respect to a global team reward optimizes the factorization of the jointed -Value.
@@
Issues found when using Q-Learning in an end-to-end GPU setting. As discussed in the paper’s results section, PPO demonstrates a clear advantage over Q-Learning for our benchmarked environments, both in agent performance and training runtime. The speed differential is caused by the optimal sampling/replay ratio for Q-Learning methods becoming rapidly unbalanced as the number of parallel environments increases, which requires us to use fewer parallel environments than we use with PPO. PPO also has a major advantage over Q-Learning in that it does not use a replay buffer, which can occupy a significant amount of GPU memory. Secondly, our experiments empirically showed PPO to be more stable during training.
A possible workaround is to increase the replay ratio by performing multiple update steps per training episode, which nevertheless affects computational efficiency. A better solution is to implement a distributed framework, separating the learning and sampling process, which is also out-of-scope for this work.
Appendix E Speed Comparison
The runs reported in Figures 3 and 5(c) were all run on the same system featuring two NVIDIA GeForce RTX 4090s (although only one was used for training), an Intel(R) Xeon(R) Silver 4316 CPU (20 cores with 40 threads), and 132 GB of RAM. We report the average environment steps per second over the entire RL training process, which for JaxMARL includes any compilation time. For Table 3, all results were collected on a single NVIDIA A100 GPU and AMD EPYC 7763 64-core processor. Environments were rolled out for 1000 sequential steps.
In Figure 11 we repeat the analysis, reported in the main paper for QMIX, of JaxMARL’s ability to train multiple seeds in parallel for IPPO. Training this way allows training agents many thousands of times faster, with a 12500x speed up in the MPE simple spread environment.
Appendix F Training & Correctness Results
F.1 Overcooked
We train IPPO, VDN and IQL agents in Overcooked and present their aggregate performance in Figure 12(a). IPPO performs better than the Q-Learning methods in inter-quartile mean and mean, in line with our more general findings. During training, we use the same shaped reward as stated in the original Overooked paper [10], which is added to the score of the game with a factor that is decayed from 1 to 0 during the first half of training. We don’t train MAPPO and QMIX for this task because, in Overcooked, agents can observe the entire state of the map. Therefore, there is no partial observability that can be improved through centralized training. We demonstrate correspondence by training an IPPO policy with JaxMARL on our implementation and evaluating the policy over 10 rollouts for both our Overcooked implementation and the original. Results are shown in Figure 14 with the similarity in performance demonstrating their equivalence.
F.2 MABrax
@@
The performance of IPPO on ant_4x2, humanoid_9|8, hopper_3x1 and walker2d_2x3 is reported in Figure 15, with hyperparameters reported in Table 7.
F.3 MPE
@@
Performance of -Learning baselines in all the MPE scenarios are reported in Figure 17 and Figure 18. The upper row represents cooperative scenarios, with results for all our -learning baselines reported. The bottom row refers to competitive scenarios, and results for IQL are divided by agent types. Hyperparameters are given in Table 8.
F.4 SMAX
The performance of different algorithms in SMAX versus MAPPO in SMAC is shown in Figure 19. Hyperparameters for IPPO and the -learning methods are given in Table 6 and Table 8 respectively. Some maps are significantly more difficult in SMAX, such as 10m_vs_11m, whereas some are much easier such as 3s_vs_5z.
F.5 Hanabi
Appendix G Hyperparameters
| Parameter | MPE | SMAX | Hanabi | Overcooked |
| PPO | ||||
| # training timesteps | ||||
| # parallel environments | 16 | 64 | 1024 | 64 |
| # rollout steps | 128 | 128 | 256 | |
| Adam learning rate | ||||
| Anneal learning rate | True | True | True | True |
| Update epochs | 2 | 4 | 4 | |
| Minibatches per epoch | 2 | 4 | 16 | |
| 0.99 | 0.99 | 0.99 | ||
| 0.95 | 0.95 | 0.95 | ||
| Clip range | 0.2 | 0.2 | 0.2 | |
| Entropy coefficient | 0.0 | 0.01 | 0.01 | |
| Value loss coefficient | 0.5 | 0.5 | 0.5 | |
| Maximum gradient norm | 0.5 | 0.5 | 0.5 | |
| Activation | tanh | relu | relu | relu |
| Feed-forward network | ||||
| Number of layers | 2 | - | 2 | 2 |
| Fully-connected width | 64 | - | 512 | 64 |
| Recurrent network | ||||
| Hidden width | 128 | 128 | 128 | - |
| Number of full-connected layers | 2 | 2 | 2 | - |
| Fully-connected width | 128 | 128 | 128 | - |
| Parameter | Ant | HalfCheetah | Walker |
| PPO | |||
| # training timesteps | |||
| # parallel environments | 64 | ||
| # rollout steps | 300 | ||
| Adam learning rate | |||
| Anneal learning rate | True | ||
| Update Epochs | 4 | ||
| Minibatches per epoch | 4 | ||
| 0.99 | |||
| 1.0 | |||
| Clip range | 0.2 | ||
| Entropy coefficient | |||
| Value loss coefficient | 4.5 | 0.14 | 1.9 |
| Maximum gradient norm | 0.5 | ||
| Activation | tanh | ||
| Feed-forward network | |||
| Number of layers | 2 | ||
| Fully-connected width | 64 |
| Parameter | MPE | SMAX | Overcooked |
| Q-Learning | |||
| # training timesteps | |||
| # parallel environments | 8 | 16 | 32 |
| # rollout steps | 26 | 128 | 1 |
| Adam learning rate | |||
| Anneal learning rate | True | False | True |
| Buffer size | |||
| Buffer batch size | 32 | 32 | 128 |
| start | 1.0 | 1.0 | 1.0 |
| finish | 0.05 | 0.05 | 0.05 |
| decay | 0.1 | 0.1 | 0.2 |
| Hidden size | 64 | 512 | 64 |
| Maximum gradient norm | 25 | 10 | 1 |
| 1.0 | 1.0 | 1.0 | |
| Update epochs | 1 | 8 | 4 |
| Learning starts at timestep | |||
| 0.9 | 0.99 | 0.99 | |
| QMIX specific | |||
| Mixed embedding width | 32 | 64 | - |
| Mixer hypernet width | 128 | 256 | - |
| Mixer initial scale | - |